Circular No. 6617

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 970402 L. Piro, M. Feroci, and E. Costa, Istituto di Astrofisica Spaziale, CNR,

Frascati; L. A. Antonelli, F. Fiore, and P. Giommi, BeppoSAX, Rome; A. Owens and A. Parmar, Space Science Department, ESTEC; S. Molendi, Istituto di Fisica Cosmica e Tecnologie Relative, CNR, Milano; G. Cusumano, Istituto di Fisica Cosmica e Applicazioni Informatica, CNR, Palermo; J. in ’t Zand, J. Heise, and R. Jager, Space Research Organization of the Netherlands, Utrecht; F. Frontera and G. Zavattini, Università di Ferrará; L. Nicastro and E. Palazzi, Istituto Tecnologie e Studio Radiazioni Extraterrestri, CNR, Bologna; L. Salotti, OCC, Nuova Telespazio, Rome; L. Di Ciolo, G. D’Andreta, R. Ricci, G. Spoliti, A. Coletta, M. Capalbi, and M. R. Daniele, BeppoSAX, Rome; and R. C. Butler, Agenzia Spaziale Italiana, Rome, report: “The BeppoSAX GRBM/WFC error box of GRB 970402 (IAUC 6610) was observed with the BeppoSAX Narrow Field Instruments on Apr. 3.280 UT for 25 000 s (8 hr after the γ-ray burst). A previously unknown x-ray source, SAX J1450.1–6920, has been detected by the MECS and LECS at the same position (α = 14^{h}50^{m}06^{s}, δ = −69^{o}20′.0, equinox 2000.0; estimated error radius 50′′). This position lies at the center of the reported BeppoSAX WFC error box (IAUC 6610). The source countrate was (3.0 ± 0.7) × 10^{−3} counts s^{−1}, corresponding to a flux of (1.5 ± 0.5) × 10^{−13} erg cm^{−2} s^{−1} in the MECS (2–10 keV), and (2.0 ± 0.6) × 10^{−3} counts s^{−1} in the LECS (0.5–5 keV), corresponding to (2.0 ± 0.6) × 10^{−13} erg cm^{−2} s^{−1} in the range 0.5–5 keV. The flux conversion has been derived assuming the galactic column density and a power law with photon index 2. The source showed a decreasing trend during the observation. The field was observed again on Apr. 4.634 for 50 000 s. No source was detected in the position of SAX J1450.1–6920. The 3σ upper limit on the source flux implies a decrease by a factor of 3 or more. Given the similarity with the behavior of the x-ray transient associated with GRB 970228 (IAUC 6576), we conclude that BeppoSAX has detected the x-ray afterglow of GRB 970402.”

SUPERNOVA 1997bp IN NGC 4680 G. V. Williams, Smithsonian Astrophysical Observatory, provides the

following measurement of SN 1997bp from the image by N. Grogin (cf. IAUC 6613): α = 12^{h}46^{m}53^{s}.75, δ = −11^{o}38′33′′.2 (equinox 2000.0), which is 14′′.1 west and 19′′.8 south of the nucleus of NGC 4680.

1997 April 8  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green